### v2

# Toward Universal Embeddings in Digital Pathology through Multimodal LLMs

Abstract
Pathology plays a critical role in diagnosing a wide range of diseases, yet existing approaches often rely heavily on task-specific models trained on extensive, well-labeled datasets. These methods face sustainability challenges due to the diversity of pathologies and the labor-intensive nature of data collection. To address these limitations, we highlight the need for universal multimodal embeddings that can support multiple downstream tasks. Previous approaches involve fine-tuning CLIP-based models, which handle images and texts separately, limiting their ability to capture complex multimodal relationships. Additionally, these models are evaluated across diverse datasets without a unified benchmark. In this paper, we explore the possibility of applying Multimodal Large Language Models (MLLMs) to generate pathology universal embeddings to address these challenges. Our contributions can be summarized in the following aspects: 1) We propose MLLM4PUE, a novel framework that leverages MLLMs to generate embeddings for various pathology downstream tasks. 2) We further introduce the Pathology Multimodal Embedding Benchmark (PMEB), a comprehensive benchmark designed to assess the quality of pathology multimodal embeddings, which comprises 16 original tasks drawn from 15 datasets. 3) Extensive experimental results demonstrate the superiority of MLLM4PUE, illustrating MLLM-based models can effectively support a wide range of downstream tasks and unify the research direction for foundation models in pathology.
1 Introduction
Figure 1: Performance comparison across benchmark datasets, showing that our proposed MLLM4PUE achieves baseline methods across most of tasks. Weight F1 (wF1) is reported for classification task. For the retrieval and composed retrieval task, we report recall@5 scores. This highlights the advanced multimodal fusion capabilities of MLLM4PUE.
Pathology remains the gold standard for diagnosing a wide range of diseases [41, 32]. With advancements in artificial intelligence (AI), digital pathology has emerged as a promising field, using AI to address challenges across various tasks such as cancer diagnosis [29, 38], metastasis detection [43, 21], mutation prediction [8, 33], and survival analysis [40, 39, 44]. These achievements, however, rely on models trained on large, well-labeled datasets. Given the hundreds of tumor types cataloged in the WHO classification111tumourclassification.iarc.who.int/ and the labor-intensive nature of data collection and annotation [11], developing separate models for each pathology task is impractical, especially for rare diseases with limited data [23, 30]. Furthermore, while many efforts focus on image, pathology often involves text, including diagnostic reports and research papers [6, 7, 31, 13]. Thus, there is a growing need to shift toward task-agnostic models that generate universal multimodal embeddings capable of supporting different pathology tasks.
Due to the success of CLIP [34] model, several studies [30, 15, 37, 16, 31] have adapted CLIP-based models for pathology, training on extensive image-text pairs to extract embeddings that enable zero-shot evaluation across various pathology tasks. However, these methods encounter two significant challenges. First, CLIP [34] processes images and text independently, limiting its ability to capture complex interactions between modalities [18]. Additionally, CLIP [34] can only handle text inputs with a maximum length of 74 tokens, which restricts its ability to leverage longer and more detailed textual descriptions. Second, evaluations of these models have mostly centered on isolated tasks, such as classification or retrieval [30, 37, 16], which separate image and text components and do not fully assess multimodal integration capabilities. Moreover, inconsistencies in datasets and sample selection criteria across studies lead to non-reproducible and incomparable results. These challenges emphasize the need for a more integrated model and a standardized benchmarking framework to advance multimodal embeddings in pathology.
To tackle the first challenge, we propose MLLM4PUE, a framework that uses a MLLM to generate universal multimodal embeddings for pathology. Unlike CLIP [34], which processes images and text separately, MLLM employs a transformer-based architecture to fully integrate these modalities, allowing it to learn from complex image-text relationships [28, 27]. Moreover, MLLM can handle longer textual inputs compared to CLIP, overcoming the 74-token limitation and enabling richer contextual understanding, which is important in pathology area. To harness the capability of MLLM in generating embeddings, we take the following steps. First, we employ summarization prompts to guide the MLLM to output embeddings. Next, we fine-tune MLLM with a LoRA module using pathology image-text pairs, adapting it specifically for the pathology domain.
To address the second challenge, we introduce the Pathology Multimodal Embedding Benchmark (PMEB) to establish a standardized evaluation platform for multimodal embeddings in pathology. This benchmark encompasses three meta-task categories: retrieval, composed retrieval, and classification. To ensure the tasks effectively assess multimodal embeddings, we have reformulated them to better suit this purpose. Each task provides the model with prompts to process a query and identify the correct target from a set of candidates. These queries and targets can consist of images, text, or a combination of both, enabling a flexible and comprehensive evaluation of multimodal capabilities. PMEB offers a standardized and reproducible approach for assessing multimodal embeddings in pathology. Overall, our contributions are summarized as follows:
  • We introduce MLLM4PUE, a framework converting pathology images and texts into robust embeddings within a unified model. This is the first time to explore the possibility of re-purposing MLLM for extracting embeddings in pathology area.
  • We provide PMEB, a comprehensive benchmark for evaluating multimodal embeddings in pathology, encompassing a broad range of tasks to support robust and generalizable assessment of these embeddings.
  • MLLM4PUE effectively represents multimodal information, consistently surpassing the performance of baseline models across amounts of evaluated tasks, as demonstrated by the quantitative results in Fig. 1.
  • Our approach offers new insights into digital pathology by adopting MLLM for universal embeddings. Future research can build on this work to unify the development of models for diverse pathology tasks.
Meta-Task Data source Query Target Original Tasks Original samples Selected samples
Retrieval Arch-book [12] I T Image-text retrieval 1,306 1,306
Arch-book [12] T I Text-image retrieval 1,306 1,306
Arch-pubmed [12] I T Image-text retrieval 1,923 1,923
Arch-pubmed [12] T I Text-image retrieval 1,923 1,923
Composed Quilt-VQA [35] I+T T Histopathology VQA 985 724
retrieval Quilt-VQA-RED [35] I+T T Histopathology VQA 335 252
Classification Bach [2] I T Breast tissue 399 399
CRC-100k [19] I T Colorectal cancer 100,000 1,000
Databiox [19] I T Colorectal cancer 922 922
Digestpath [9] I T Signet ring cell 455 455
Digestpath [9] I T Colonoscopy tissue 660 660
Kathercolon [20] I T Colorectal cancer tissue 7,180 1,000
LC25000 [4] I T Colon adenocarcinoma 10,000 1,000
LC25000 [4] I T Lung adenocarcinoma 15,000 1,000
Osteo [3] I T Osteosarcoma 1,144 1,144
Renalcell [5] I T Renal tissue 36,687 1,000
Sicap [36] I T Gleason pattern 12,081 1,000
Skincancer [22] I T Skin cancer tissue 129,369 1,000
Wsss4luad [14] I T LUAD tissue 10,091 1,000
Table 1: The statistics of our PMEB benchmark: 15 original tasks from 14 datasets. I and T denote image and text, respectively.
2 Related Work
Vision-language Models in Digital Pathology. In digital pathology, recent studies have adopted the CLIP [34] model, leveraging paired image-text data within a contrastive learning framework to align similar image-text embeddings while separating dissimilar ones [30, 15, 37, 16, 17]. For instance, PLIP [15] fine-tunes CLIP on large collections of image-text pairs sourced from platforms like Twitter and other public datasets. Similarly, PathCLIP [37] and QuiltNet [16] scale up pathology-specific data for training. CONCH [31] employs over 1.17 million image-caption pairs to fine-tune CoCa [42] model. Despite these advancements, these methods fine-tune either CLIP [34] or CoCa [42] using pathology image-caption pairs and rely on separate encoders for processing images and text. In contrast, we propose an MLLM-based framework to capture universal multimodal embeddings for pathology by integrating image and text data within a unified model. This integrated framework enhances performance across a wide range of digital pathology tasks, offering a more holistic and effective solution.
Multimodal Embedding Learning. Creating effective multimodal embeddings remains a challenging area of research. CLIP  [34] addresses this by employing separate encoders for images and text, aligning them in a common space through contrastive learning. This approach has influenced many subsequent models, such as BLIP [25] and CoCa [42]. While these models excel in a variety of tasks, the use of distinct encoders restricts their capacity to fully integrate visual and textual data, which is a key requirement for tasks that involve combined visual-language inputs, such as search queries involving both images and text.
With the rise of large language models (LLMs), MLLMs have extended LLMs to process multiple data modalities, achieving notable progress in understanding and reasoning across diverse input types [28, 27, 24]. Although MLLMs exhibit strong capabilities in interpreting multimodal content and following complex instructions, there is still limited research on their application in creating embeddings. The recent work E5-V [18] fine-tunes an LLM with summarization prompts and text-only data to extract embeddings, then integrates a vision module to obtain multimodal embeddings for zero-shot multimodal retrieval tasks. However, for pathological adaptation, E5-V [18] requires a domain-specific LLM pretrained on pathology instruction data, which demands detailed and standardized annotations. In contrast, our framework, MLLM4PUE can fine-tune MLLM with a LoRA module using existing pathology image-text pairs, bypassing these prohibitive requirements.
Figure 2: Overview of the MLLM4PUE framework. (a) Illustration of pathology image-text contrastive learning with our proposed method. With differents prompts, MLLM can generate embeddings for multimodal inputs. Image and text embeddings are aligned using a contrastive loss. (b) Zero transfer for classification. (c) Zero transfer for composed retrieval.
3 Methodology
In this section, we first introduce the Pathology Multimodal Embedding Benchmark (PMEB), a comprehensive benchmark for evaluating multimodal embeddings in diverse pathology tasks. Then, we present the MLLM4PUE framework, which integrates multimodal embeddings, contrastive learning, and zero-shot transfer methods to align and adapt pathology image-text data effectively, as illustrated in Fig. 2.
3.1 Pathology Multimodal Embedding Benchmark
We propose PMEB, a comprehensive benchmark designed to evaluate pathology multimodal embeddings. PMEB include cancer tissue classification [2, 4, 22], Gleason pattern grading [36], and cell identification [9], among other tasks. As shown in Table 1, PMEB comprises 15 original tasks from 14 datasets, organized into three meta-tasks: retrieval, classification, and composed retrieval. All tasks are reformulated to assess multimodal embeddings, where the model receives an instructional prompt and a query, which may consist of text, images, or both, and selects the correct target from a set of options.
In the retrieval task, the query can be either an image or text, with the target being the corresponding paired text or image. For the classification task, queries are images with targets as text labels representing various classes. The composed retrieval task involves queries that combine an image with text formatted as a question, with the target being the expected answer. This composed retrieval task, specifically designed for multimodal embedding evaluation, incorporates integrated visual and language inputs and has not been applied in previous methods.
Due to the large size of some datasets, we standardize their sizes. For datasets containing more than 1,000 samples (except Osteo [3]), we choose a random sample of 1,000, ensuring the original category distribution is preserved. For the composed retrieval task, we select only open-ended questions from the data source, where answers are more complex than a simple “yes” or “no”. Additionally, since there is an overlap between questions and answers in the VQA data source, we utilize ChatGPT [1] for data cleaning to ensure clarity and accuracy in task evaluation. Further details of each dataset within PMEB are provided in Supplementary Materials.
3.2 Multimodal Embeddings with MLLM
We first introduce how our proposed MLLM4PUE leverages MLLMs for universal multimodal embedding learning. Unlike CLIP, which directly generates embeddings for multimodal inputs, MLLMs are inherently generative models and require a different approach to produce embeddings. We uses a prompt-based strategy specifically designed to guide MLLMs in embedding multimodal data.
[…]
The prompt design incorporates two components: first, the word Summarize directs the MLLM to distill the information of the multimodal input. The second part, in one word:, instructs the MLLM to compress this distilled information into a single token, thereby facilitating a unified multimodal embedding. Once the extended input $q_{input}$ is constructed, it is processed by the MLLM $f_{\varphi}$ to generate embeddings: $h=f_{\varphi}(q_{input})$, where $h$ is the embedding of last token obtained from the MLLM output, as illustrated in Fig. 3. This approach tailors the generation capabilities of MLLMs to effectively produce and utilize multimodal embeddings, expanding their application potential beyond traditional generative outputs.
3.3 Pathology Contrastive Learning
Next, we describe the approach for projecting pathological data into a latent embedding space using conservative learning. As illustrated in Fig. 2, given a batch of $N$ paired image and text samples $\{(v_{n},t_{n})\}_{n=1,\dots,N}$ and MLLM $f_{\varphi}$, for each sample $(v_{i},t_{i})$, we feed them into $f_{\varphi}$ to get the image and text embeddings $(h_{vi},h_{ti})$ as described in Sec. 3.2.
The objective is to build a latent space for pathology image-text embeddings that maximizes the similarity between embeddings of the paired samples in the batch, while minimizing the similarity between embeddings of the $N-1$ incorrect pairs. We optimize an infoNCE contrastive Loss $\mathcal{L}$ to train our model.
$\mathcal{L}=-(\log\frac{e^{cos(h_{vi},h_{ti})/\tau}}{\sum_{j=1}^{n}e^{cos(h_{vi},h_{tj})/\tau}}+\log\frac{e^{cos(h_{ti},h_{vi})/\tau}}{\sum_{j=1}^{n}e^{cos(h_{ti},h_{vj})/\tau}}),$ (1)
where $\tau$ is a temperature parameter and $cos(h_{vi},h_{ti})$ and $cos(h_{ti},h_{vi})$ represent the cosine similarity in both directions in contrastive learning.
Figure 3: Illustration of the MLLM4PUE framework for generating multimodal embeddings using an MLLM. The image is processed through an Image Encoder and projection layer, with a prompt guiding the MLLM to distill the image content into a single-word embedding. The final embedding $h$ is extracted from the last token as the unified multimodal representation.
3.4 Zero-shot Transfer Evaluation
Our model, trained primarily on retrieval tasks focused on image-text pairs, requires adaptation to handle zero-shot classification and composed retrieval tasks. The details of how we modify the model for these different zero-shot downstream tasks are introduced as follows.
For zero-shot classification, we adopt a prompt-based method inspired by the CLIP model [34]. In this approach, each class name is expanded into a sentence using a specific template. For instance, the class name “Colon adenocarcinoma” is converted into the sentence “An H&E image of Colon adenocarcinoma” using the template “An H&E image of { }.” We apply this method to create sentences for all class names. Our model then computes embeddings for these sentences and the test images, extracting the last token of the MLLM output as described in Section 3.2. The similarity between these embeddings is calculated as outlined in Section 3.3, and test image labels are assigned based on the highest similarity scores.
[…]
Task Model Arch-PubMed [12] Arch-book [12]
Recall@5 Recall@10 Recall@50 Recall@5 Recall@10 Recall@50
i2t E5-V [18] 0.006 0.011 0.052 0.009 0.018 0.067
PLIP [15] 0.037 0.067 0.185 0.096 0.152 0.393
MLLM4PUE-O 0.105 0.160 0.366 0.185 0.264 0.575
PathCLIP [16] 0.275 0.388 0.680 0.152 0.234 0.482
MLLM4PUE-P 0.372 0.495 0.782 0.192 0.283 0.603
QuiltNet [16] 0.069 0.111 0.273 0.116 0.168 0.384
MLLM4PUE-Q 0.122 0.177 0.407 0.182 0.248 0.502
t2i E5-V [18] 0.006 0.011 0.052 0.009 0.018 0.067
PLIP [15] 0.037 0.067 0.181 0.112 0.164 0.419
MLLM4PUE-O 0.114 0.173 0.385 0.193 0.280 0.600
PathCLIP [16] 0.236 0.348 0.630 0.137 0.196 0.445
MLLM4PUE-P 0.297 0.399 0.688 0.185 0.277 0.555
QuiltNet [16] 0.056 0.092 0.237 0.100 0.152 0.389
MLLM4PUE-Q 0.135 0.193 0.433 0.210 0.302 0.584
Table 2: Performance comparison with baseline methods on two retrieval datasets, reporting recall metrics at different thresholds. Task i2t and t2i denote image-to-text and text-to-image retrieval, respectively. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [15], PathCap [37], Quilt1M [16], respectively. Bold values indicate the best performance. Same color represents models trained on the same dataset.
4 Experiments
Training datasets. We leverage three public datasets, Openpath [15], PathCap [37] and Quilt1M [16], to construct a robust training foundation for pathology image-text pairs. The Openpath dataset [15] is assembled by gathering data from Twitter and LAION using popular pathology hashtags. Since PLIP[15] provided only Twitter IDs, we retrieve the corresponding data from Twitter and LAION using these IDs, resulting in a dataset of 138,874 image-text pairs after denoising. The PatchCap dataset [37] comprises 207K high-quality samples from authoritative sources, and we use all provided image-text samples. Similarly, the Quilt1M dataset [16] aggregates pathology image-text pairs from several public sources. As many images in Quilt1M are associated with multiple captions, we concatenate the captions for each image to create comprehensive entries. Through these careful preparations, we compile an extensive training dataset consisting of 593,838 image-text pairs.
Baseline Methods and Metrics. We evaluate the performance of our proposed MLLM4PUE against several recent baseline models across various downstream tasks. Specifically, we compare it with three pathology CLIP-based models: PLIP [15], PathCLIP [37], and QuiltNet [16], all of which are trained on the same dataset as our method. Additionally, we compare our approach with E5-V [18], a method that fine-tunes an LLM with summarization prompts and text-only data to extract embeddings, then integrates a vision module to obtain multimodal embeddings for zero-shot multimodal retrieval tasks. For zero-shot classification tasks, we also compare our method with CONCH [31], a model trained on over 1.17 million private image-caption pairs, whereas our method is trained on fewer than 0.6 million pairs.
To evaluate the performance of retrieval and composed retrieval tasks, we use the Recall@K metric, which quantifies the proportion of relevant items contained within the top K retrieved results. For classification tasks, due to the imbalanced distribution of classes, we use the weighted F1 (wF1) score to assess performance. The weighted F1 is calculated by averaging the F1 scores for each class, with each class’s score weighted by its frequency of occurrence.
Implementation Details. Our framework is implemented in Pytorch and we use LLaVA-NeXT-8B [24] as our backbone MLLM. To save the GPU memory, we use QLoRA [10] and gradient checkpointing with DeepSpeed ZeRO-2 to fine-tune the MLLM, and all the images are resized to 336 × 336 pixels. The temperature is set to 0.02. All experiments are run on six H100 GPUs with a gradient accumulation batch size of 144 and a learning rate of 4e-4.
Model Bach
Data
biox
Digest
Digest
CRC
100k
Kather
7k
LC
colon
LC
lung
Osteo
Renal
cell
Sicap
Skin
cancer
Wsss
4luad
E5-V [18] 0.100 0.162 0.197 0.049 0.224 0.206 0.333 0.209 0.300 0.051 0.230 0.017 0.132
CONCH [31] 0.511 0.413 0.835 0.824 0.680 0.625 0.890 0.650 0.382 0.393 0.359 0.593 0.772
PLIP [15] 0.124 0.194 0.696 0.252 0.503 0.504 0.697 0.743 0.554 0.413 0.080 0.311 0.394
MLLM4PUE-O 0.561 0.437 0.784 0.480 0.642 0.509 0.958 0.744 0.581 0.507 0.284 0.411 0.636
PathCLIP [37] 0.528 0.282 0.785 0.707 0.611 0.492 0.937 0.887 0.668 0.359 0.267 0.332 0.648
MLLM4PUE-P 0.577 0.388 0.799 0.900 0.763 0.783 0.937 0.886 0.751 0.543 0.502 0.504 0.754
QuiltNet [16] 0.403 0.318 0.491 0.564 0.564 0.591 0.884 0.801 0.578 0.435 0.322 0.467 0.364
MLLM4PUE-Q 0.437 0.434 0.868 0.851 0.686 0.668 0.886 0.866 0.613 0.516 0.440 0.507 0.758
E5-V [18] 0.251 0.324 0.367 0.169 0.299 0.306 0.500 0.348 0.385 0.166 0.337 0.053 0.267
CONCH [31] 0.559 0.444 0.832 0.802 0.703 0.644 0.890 0.665 0.439 0.447 0.405 0.607 0.783
PLIP [15] 0.251 0.306 0.708 0.253 0.532 0.541 0.721 0.749 0.551 0.488 0.179 0.290 0.450
MLLM4PUE-O 0.569 0.448 0.793 0.429 0.666 0.546 0.958 0.775 0.624 0.518 0.426 0.453 0.671
PathCLIP [37] 0.556 0.386 0.784 0.666 0.610 0.489 0.937 0.888 0.698 0.350 0.362 0.378 0.688
MLLM4PUE-P 0.599 0.386 0.812 0.892 0.767 0.784 0.937 0.886 0.744 0.576 0.509 0.496 0.760
QuiltNet [16] 0.441 0.413 0.628 0.503 0.607 0.586 0.904 0.802 0.588 0.434 0.338 0.458 0.449
MLLM4PUE-Q 0.479 0.441 0.866 0.849 0.695 0.661 0.887 0.872 0.609 0.494 0.444 0.520 0.761
Table 3: Performance comparison with baseline methods on zero-shot classification tasks, reporting weighted F1 (wF1) (top) and accuracy (bottom) metrics. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [15], PathCap [37], Quilt1M [16], respectively. Bold values indicate the best performance. Same color represents models trained on the same dataset.
5 Results
5.1 Zero-shot Cross-modal Retrieval
We first evaluate our methods on the retrieval task on Arch-PubMed [12] and Arch-book datasets [12], where the query is an image or text, with the target being the corresponding paired text or image. Recall@k metric is employed to assess the model’s ability to retrieve relevant pathology image-text pairs among top k results. As shown in Table 2, our models consistently outperform the baseline models, especially at higher recall values. Notably, in the image-to-text (i2t) task on Arch-book dataset [12], our MLLM4PUE trained on Openpath dataset [15] achieves a Recall@5 of 0.185, indicating an approximate 9% improvement over PLIP [15]. Moreover, our PathCap-pretrained model surpasses the performance of PathCLIP [37] on both benchmark datasets. A similar enhancement is observed in the text-to-image (t2i) retrieval task. For instance, on the Arch-PubMed dataset [12], our model achieves a Recall@10 of 0.193, marking a 10% increase in performance. These results collectively demonstrate that our framework excels at capturing multimodal embeddings more effectively than CLIP-based models with the same training data. Furthermore, it is important to note that the E5-V model [18] yields significantly lower recall scores across all tasks and datasets due to the lack of fine-tuning with pathology-specific data.
Model Quilt-VQA [35] Quilt-VQA-RED [35]
Recall@5 Recall@10 Recall@50 Recall@5 Recall@10 Recall@50
E5-V [18] 0.214 0.323 0.495 0.351 0.441 0.668
PLIP [15] 0.191 0.224 0.349 0.250 0.286 0.500
MLLM4PUE-O 0.239 0.334 0.583 0.373 0.460 0.790
PathCLIP [16] 0.192 0.242 0.413 0.262 0.337 0.603
MLLM4PUE-P 0.218 0.329 0.541 0.363 0.468 0.790
QuiltNet [16] 0.180 0.209 0.336 0.222 0.274 0.460
MLLM4PUE-Q 0.475 0.598 0.823 0.675 0.754 0.948
Table 4: Zero-shot composed retrieval performance on the Quilt-VQA [35] and Quilt-VQA-RED [35] datasets, with Recall metrics reported at various thresholds. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [15], PathCap [37], Quilt1M [16], respectively. Bold values indicate the best performance. Same color represents models trained on the same dataset.
5.2 Zero-shot Patch Classification
We then evaluate the performance of zero-shot classification on patch-level pathology images across 13 classification tasks from our PMEB. In this downstream task, each query is an image, and the target is an extended sentence of the class name as described in Section 3.4. Table 3 illustrates that our MLLM4PUE consistently surpasses baseline models on most datasets. Notably, with the same training data, MLLM4PUE-O outperforms PLIP [15] across all datasets. Similarly, MLLM4PUE-Q achieves the best performance compared to QuiltNet [16], demonstrating the effectiveness of our pretraining strategy on the same training dataset. While CONCH [31] attains the highest score on SkinCancer [22] and Wsss4luad [14], its superior performance stems from utilizing a much larger private dataset (1.16M). In contrast, our model, MLLM4PUE-P, outperforms all baseline models across seven datasets with only 0.2 million training samples, and is only 0.02 less than CONCH [31] on the Wsss4luad dataset [14]. This underscores the advantages of leveraging vision-language alignment in pathology representation learning. These findings collectively highlight the effectiveness of our approach, especially considering that our models were trained on fewer samples, further demonstrating their data efficiency in pathology-specific zero-shot classification. The accuracy metric shows a similar trend, where MLLM4PUE-P outperforms the baselines on most datasets.
5.3 Zero-shot Composed Retrieval
[…]
5.4 Visualization of Modality Gap
To illustrate the modality gap, we employed t-SNE visualization to compare the embeddings of our MLLM4PUE-P model with those of PathCLIP [37], both trained on the Patchcap dataset [37]. As shown in Fig 4, the qualitative analysis clearly shows that the embeddings produced by MLLM4PUE-P form tighter clusters in the low-dimensional space, whereas those from PathCLIP [37] are more dispersed. This difference reflects a significant improvement in cross-modal alignment of our method.
Figure 4: Distribution of image and text embeddings on Arch-PubMed [35] and Arch-Book [35] dataset.
To further quantify the modality gap, we adopted the metric $||\Delta||_{gap}$ as proposed in Liang et al.  [26]. Our experiments reveal that, on the Arch-PubMed dataset [35], our method reduces the gap from 0.9284 to 0.2587, and on the Arch-Book dataset [35], from 1.0408 to 0.3739. These results provide strong evidence that our model can significantly narrow the gap between image and text embeddings.
5.5 Ablation Studies
Model Quilt-VQA [12] Quilt-VQA-RED [12]
Recall@5 Recall@10 Recall@50 Recall@5 Recall@10 Recall@50
QuiltNet [16] only question 0.170 0.189 0.275 0.187 0.234 0.440
add embedding 0.180 0.209 0.336 0.222 0.274 0.460
MLLM4PUE only question 0.291 0.362 0.539 0.429 0.488 0.694
add embedding 0.394 0.500 0.746 0.520 0.619 0.909
prompt 0.475 0.598 0.823 0.675 0.754 0.948
Table 5: Ablation study of image and question embedding fusion methods on composed retrieval task. All models are trained on the Quilt1M [16] dataset. Recall metrics are reported. Bold values indicate the best performance.
Robustness of PMEB. Since some datasets in our PMEB are super large, we choose a random sample of 1,000. To assess the robustness of our benchmark, we conduct an ablation study by computing the wf1 scores using 1,000, 3,000, 5,000, and all available samples for each dataset. As shown in Figure 5, the performance change between 1,000 and the full dataset remains within a small margin across all datasets, confirming that using only 1,000 samples provides a reliable approximation of the complete dataset’s performance. Given that the performance difference is typically within 0.01 to 0.04 across different sample sizes, using 1,000 samples significantly reduces computational costs while maintaining reliable performance. This makes PMEB highly efficient for large-scale experiments without sacrificing meaningful evaluation. We also report the result of accuracy metric and the results in CONCH [31] and PathCLIP [37] model in the Supplementary Materials, which shows the same trend.
Figure 5: Ablation study on the sample numbers of our benchmark on our MLLM4PUE-P model. We computed the wF1 scores using 1,000, 3,000, 5,000, and all available samples for each dataset.
Embedding Fusion Methods. This experiment is designed to validate the quality of our benchmark and highlight the effectiveness of our MLLM4PUE in modality fusion compared to CLIP-based model in composed retrieval tasks. As shown in Table 5, MLLM4PUE outperforms QuiltNet [16] in every setting. First, comparing “only question” results, our MLLM4PUE shows significant improvement compared with QuiltNet [16], while both models demonstrate substantial performance gains when incorporating visual information via “add embedding”. This consistent pattern across both models confirms that our benchmark has been properly curated, requiring true integration of visual and textual information rather than allowing models to infer answers solely from questions.
Second, the results reveal that QuiltNet [16], a CLIP-based model, shows little difference between using only the question text and integrating image and text embeddings. This indicates that CLIP’s method lacks robust modality integration, as the inclusion of visual information fails to significantly enhance its effectiveness. Conversely, our model MLLM4PUE demonstrates substantial improvements by effectively fusing image and text information. The true strength of our approach becomes evident when employing a prompt-based method, leading to significantly higher recall scores. Specifically, it achieves Recall@5 values of 0.475 on Quilt-VQA and 0.675 on Quilt-VQA-RED, representing approximately an 8% and 12% improvement over the simple addition method, respectively, and an 18% and 25% improvement over using the question text alone. These findings clearly indicate that our model’s ability to integrate both visual and textual modalities results in significant enhancements in retrieval accuracy. This surpasses both mere modality addition and the question-only configuration, highlighting the effectiveness of our modality fusion technique for composed retrieval tasks. Our approach provides advantages that cannot be achieved through separate handling or limited integration of text and image data.
6 Conclusion
In this paper, we have explored a novel paradigm for applying MLMMs to generate multimodal universal embeddings for a broad range of pathology downstream tasks. By applying summarization prompts and inserting lightweight LoRA modules, our proposed framework, MLLM4PUE, can effectively integrates image and text data, and addresses the limitations of previous CLIP-based methods. We also introduce PMEB, which provides a standardized and comprehensive benchmark for evaluating multimodal capabilities, promoting reproducibility and comparability in the field. Comprehensive experiment results across 15 diverse zero-shot retrieval and classification tasks underscore the versatility and effectiveness of our method. We anticipate that our framework will provide valuable insights into the pathology vision-language foundation models and inspire further research in this area.